### v1

@@ 6 Conclusion
    Despite the wide range of flux states observed, the source exhibits only moderate HR variability across both intra-day and long-term timescales.
@@
    The RMS-flux relation is recognized as a significant characteristic of AGN and micro-quasars, potentially linking accretion disk dynamics to jet properties. Our investigation of this relation in the source does not reveal a straightforward correlation. Complex turbulence patterns extending along the curved jet structure may obscure potential signatures of disk-jet coupling. connection.
@@
Acknowledgements.
This work was partially supported by a program of the Polish Ministry of Science under the title ‘Regional Excellence Initiative’, project no. RID/SP/0050/2024/1. TPA acknowledges the support of the National Natural Science –Foundation of China ( grant nos. 12222304, 12192220, and 12192221). MM acknowledges the International Macquarie Research Excellence Scholarship (iMQRES) program for the financial support during the research, and the ARC Centre of Excellence for All Sky Astrophysics in 3 Dimensions (ASTRO 3D), through project CE170100013.
Appendix A Light curves, hardness ratio plots, spectral fits and discrete Fourier periodograms of XMM- Newton observations of PKS 2155-304.
Figure 10: Lightcurves of Blazar PKS2155-304 from 2000 to 2014
Figure 11: Lightcurves of Blazar PKS2155-304 from 2000 to 2014
Figure 12: Lightcurves of Blazar PKS2155-304 from 2000 to 2014
Figure 13: Lightcurves of Blazar PKS2155-304 from 2000 to 2014
Figure 14: The spectral fitting of observations of PKS 2155–304 in 0.3–10 keV. Each spectra is fitted using the powerlaw, BPL and log parabolic model and the data-to-model ratio is shown in the three subpanels for each spectra in olive, red and blue colors, respectively.
Figure 15: Continue…
Figure 16: Continue…
Figure 17: Flux histograms for PKS 2155-304 observations.
Figure 18: Continue…

### v2

@@ 6 Conclusion
    The RMS-flux relation is recognized as a significant characteristic of AGN and micro-quasars, potentially linking accretion disk dynamics to jet properties. Our investigation of this relation in the source does not reveal a straightforward correlation. Complex turbulence patterns extending along the curved jet structure may obscure potential signatures of disk-jet coupling.
@@
    Despite the wide range of flux states observed, there are only two observations where clear cases of HR variability are detected on intra-day timescales.
@@
Acknowledgments
We thank the anonymous referee for their careful review and valuable suggestions, which have significantly improved the presentation of this paper. This work was partially supported by a program of the Polish Ministry of Science under the title ‘Regional Excellence Initiative’, project no. RID/SP/0050/2024/1. TPA acknowledges the support of the National Natural Science –Foundation of China ( grant nos. 12222304, 12192220, and 12192221). MM acknowledges the International Macquarie Research Excellence Scholarship (iMQRES) program for the financial support during the research, and the ARC Centre of Excellence for All Sky Astrophysics in 3 Dimensions (ASTRO 3D), through project CE170100013. RP would like to express his acknowledgement for the institutional support of the Research Centre for Theoretical Physics and Institute of Physics, Silesian University in Opava, SGS/30/2023, and for the support from the project of the Czech Science Foundation GAČR 23-07043S.